OpenStudy (anonymous):

Equation of line that goes through (1,3) (-2,-3) My options: y=-2x+1 y=2x+1 y=1/2x-1

OpenStudy (nurali):

(x-x1)/(x2-x1) = (y-y1)/(y2-y1) (x1, y1) = (1, 3) (x2, y2) = (-2, -3) Now substitute the above values in the formula: (x-1)/(-2-1)=(y-3)/(-3-3) (x-1)/-3=(y-3)/-6 -6(x-1)=-3(y-3) -6x+6=-3y+9 6x-3y+9-6=0 6x-3y+3=0 3y=6x+3 Divide by 3 y=2x+1

OpenStudy (anonymous):

Thank you(:

OpenStudy (nurali):

My Pleasure.